Find parametric equations for the segment joining the given points. (1, 3) and (3,6) c(t) = { 1 +31, 2+36 ), ostsi
Answer #1

we have

(x_1,y_1)=(1,3) \ and \ (x_2,y_2)=(3,6)

now the parametric equation of line passing through point (1, 3) and (3, 6) is,

x=x_1+(x_2-x_1)t, \ \ y=y_1+(y_2-y_1)t

x=1+(3-1)t, \ \ y=3+(6-3)t

x=1+2t, \ \ y=3+3t

hence,

c(t)=(1+2t,3+3t)